OpenStudy (anonymous):

fed up find the integral sqrt(81-x^2)/x^2 I know it's trig sub, but I get stuck at the end. I've been on this question for the past hour...

hartnn (hartnn):

what did you substitute ? "stuck at the end" ....can you show the steps what you did ? maybe i could help you spot the error :)

OpenStudy (anonymous):

x = 9sin theta dx = 9cos theta \[\sqrt(81 - 81\sin^2 \theta)/(81\sin^2\theta)9\cos \theta d \theta\] \[(1/3) (\cos^2 \theta)/(\sin^2 \theta) d \theta\] \[\cos^2 \theta = (1 + \cos2 \theta) / 2\]

OpenStudy (anonymous):

\[\sin^2 \theta = (1 - \cos2 \theta) / 2\] \[1+\cos(2\theta) / 2 * 2/1-\cos(2\theta)\]

hartnn (hartnn):

firstly, how did 1/3 come up ?? numerator and denominator, both have 81 ? secondly, you got cos^2 t/ sin^2 t which is actually cot^2 t and there's a really easy way to get integral of cot^2 t , know it ?

OpenStudy (anonymous):

The numerator sqrt(81-81sin^2) = sqrt(81(1-sin^2)) = 9sqrt(1-sin^2) and no, I don't.

hartnn (hartnn):

and dx = 9 cos t, so this 9 will also go in numerator right ?

OpenStudy (anonymous):

Yeah, I see now

hartnn (hartnn):

hint : \(\Large 1+\cot^2x = \csc^2 x\)

hartnn (hartnn):

so whats cot^2 t from there ??

OpenStudy (anonymous):

got it. thanks. =)

hartnn (hartnn):

welcome ^_^ and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)